hello there,
trying to check these questions. please help.
n^2 + 2n + 2 = (n-1)(n+3)
2b^2 -2 = 2(b-1)
2(p-q)^2 = 4p^2-q^2

Answer by stanbon(29602)

1. n^2 +2n + 2 = (n-1)(n+3)
No, that is not correct. Use FOIL on the right side and you will see
what I mean.
2. 2b^-2 = 2(b_1)
Again, that is not correct. 3b^2-2 = 2(b^2-1)= 2(b+1)(b-1)
3. 2(p-q)^2 =4p^2-q^2
Wrong again. The correct answer is 2(p^2-2pq+q^2)
Cheers,
Stan H.